Circular No. 6380

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996X IN NGC 5061 G. J. Garradd, Loomberah, Australia, and S. Nakano, Sumoto, Japan,

report the independent discoveries of an apparent supernova in NGC 5061 by Robert O. Evans (Coonabarabran, N.S.W., on Apr. 12.5 UT, mag ∼ 13, visual) and by Kesao Takamizawa (Saku-machi, Nagano-ken, on Apr. 12.618, mag 13.5, T-Max 400 film), respectively. R. and Y. Kushida (Yatsugatake) report the following precise position for SN 1996X from a CCD frame exposed on Apr. 13.480: α = 13^{h}18^{m}01^{s}.13, δ = −26^{o}50′45′′.3 (equinox 2000.0; mag 14.0), which is ∼ 52′′ west and 31′′ south of the galaxy’s center. From a CCD exposure on Apr. 13.443, Garradd provides a position for the supernova with position end figures 01^{s}.14, 46′′.0. Evans reports that no object was seen at this position on Mar. 24; Y. Kushida reports that a patrol film (Technical Pan 6415, limiting mag 15.5) exposed on 1993 Feb. 24 shows nothing at this position. Furthermore, no object is present at this location on the digital sky survey (field taken in May 1976). A nearby field star of mag ∼ 16 has position end figures 03^{s}.33, 30′′.5.

SUPERNOVA 1996W IN NGC 4027 N. B. Suntzeff reports that spectrophotometry (range 370–730 nm)

obtained by M.-T. Ruiz with the Cerro Tololo 4-m telescope on Apr. 13.1 UT confirms that this is a type-II supernova near maximum (cf. IAUC 6379), with a very blue continuum. CCD images were obtained by D. DePoy with the Cerro Tololo 0.9-m telescope, yielding the following position (using the digital sky survey) for SN 1996W: α = 11^{h}59^{m}28^{s}.98, δ = −19^{o}15′21′′.9 (equinox 2000.0), which is 17′′ west and 34′′ north of the nucleus of the galaxy. Provisional photometry gives V = 15.1 on Apr. 13.1.

COMET C/1995 O1 (HALE-BOPP) Visual m_{1} and coma-diameter estimates: Mar. 23.50 UT, 8.6, – (A.

Hale, Cloudcroft, NM, 10×50 binoculars); 27.67, 7.9, – (D. Seargent, The Entrance, NSW, Australia, 25×100 binoculars); 28.13, 8.5, 2′.6 (M. Plsek, Lelekovice, Czech Republic, 0.35-m reflector); 30.46, 8.2, – (Hale); Apr. 1.77, 7.3, ∼ 8′ (Seargent, 10×50 binoculars); 7.09, 8.3, 4′.5 (A. Baransky, Kiev, Ukraine, 0.11-m reflector).

1996 April 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green